# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented Large Language Models

### v4

@@ 5.4. Ablation Study
Figure 8. Ablation study of keywords number and Top-$k$. We compare TPIA’s ASR and STF under three conditions: without keywords, with 1-keyword, and with 2-keywords. During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
@@
Figure 9. Ablation study of FRE. We compare the differences in ASR and loss function values both with and without using FRE. 1-key1 and 1-key2 represent keeping one of the two keywords, respectively.
@@
Effects of Forward Reasoning Enhancement. We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function. Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve. It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion. To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using static instead of optimized keywords. We judge the degree of distortion based on the loss value. The results are illustrated in Figure 9, where we have three key findings. Firstly, under the same conditions, using FRE can lead to a higher ASR. Secondly, without FRE, the attack success rate is still very low ($<20\%$) even when the loss value is lower than that with FRE. Thirdly, FRE has a greater impact on attack success rate than keyword-based design. These results verify the necessity of forward reasoning enhancement in our TPIA method.
6. Discussions
6.1. Resistance to Potential Defenses
@@
In Section 3.3, we have explained why our attack can fundamentally circumvent existing malicious code detection methods. In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack. In general, we can roughly categorize these methods into detection and removal and design following potential defense methods.
@@
Detection: Semantic Analysis. Detection methods distinguish between malicious payload code containing perturbations and normal code, thereby rejecting the use of malicious payload code. Since detection methods based on code functionality are ineffective against non-functional perturbations, it is necessary to adopt a semantic analysis approach. This involves examining the differences between the benign code and the suspicious code in the embedding layer of Code LLMs. However, this method may be difficult to implement. We find that code segments injected with perturbations might not have obvious semantic differences from benign ones. Specifically, as shown in Figure 16, we construct classification prompts and extract their embedding layer features on Code LLMs, finding that the two types of code segments are almost completely indistinguishable. We present more details and analyses in Appendix A.3.
@@
Figure 10. The resistance to the removal of chars. We remove characters from the perturbation, selecting 5% for removal each time, and we calculate the average ASR for each programming language across applicable cases. As the influence of the perturbation fades, the ASR results converge to the same level as no attack baseline.
@@
Removal: Massive Deletion and Find-and-Replace. This type of defense involves removing perturbations by altering the non-functional parts of the code after introducing external code. In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect malicious code segments (as mentioned above). For example, if preserving code readability is a priority, we cannot remove all perturbations, allowing only for a lower degree of modifications. To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations. As shown in Figure 10, if the removal rate of perturbations cannot be ensured to be above 0.5, TPIA still maintains a sufficiently high ASR. On the other hand, even if readability is completely sacrificed by extensively removing comments, TPIA can still be effective through non-comment portions of the code, as indicated in Table 5. Arguably, the removal methods are likely to require significant human effort and may not yield effective results since TPIA imposes strict limitations on the length of perturbations.
@@
Table 5. ASR (Deg) for different forms of perturbations. Deg denotes the ASR degradation compared with comments form.
@@
6.2. Potential Limitations and Future Directions
@@
Although TPIA is effective in attacking the evaluated LLMs and commercial applications, we have to admit that it still has some drawbacks. Firstly, the effectiveness and stealthiness of TPIA are closely related to the keywords. We currently select keywords through a simple grid search. How to optimize keywords more efficiently is an important issue. Besides, TPIA cannot generate fully natural-looking perturbations within the context, resulting in a lack of fluency to some extent. We believe that readable and benign-looking perturbations may pose a more realistic threat and represent an important direction for future research.
7. Conclusion
@@
In this paper, we investigated an important security vulnerability of Code LLMs for code completion tasks. We revealed the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduced TPIA, a new target-specific attack paradigm without controlling the model’s training process. TPIA can circumvent mainstream detection methods (e.g., static analysis and sandboxes) by using non-functional code, instead of functional ones used in software supply chain attacks, to achieve malicious objectives. We evaluated TPIA under three popular programming languages and 31 representative cases. We showed that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation. Our work alerted a new serious security threat in using Code LLMs in practice.
Appendix A Appendix
A.1. Differences with Existing Methods
GCG attack. In the domain of generative models, there exists some adversarial prompt injection attack methods, such as universal trigger (Wallace et al., 2019) and well-known GCG (Zou et al., 2023). Taking GCG as an example, the adversarial goal of the attacker is to force LLMs to generate affirmative responses beginning with “Sure, here’s…,” thereby disrupting security alignment (Ji et al., 2024). Unlike GCG, TPIA emphasizes the feasible usage of perturbations and performance during indirect attacks. First, TPIA transfers different malicious objectives to perturbations to induce specific code generation, while GCG aims to generate the most powerful perturbation to break the safety guard. Second, TPIA considers a scenario where the users are victims and uses keyword-based perturbation to save the number of perturbation tokens, while GCG assumes that the users of LLMs are malicious attackers and allow unlimited tokens. Third, TPIA proposes a code-oriented task construction framework and forward reasoning enhancement technique to ensure effectiveness and robustness in uncertain contexts, while GCG has a controllable context and fixes the perturbation at the end of the prompt. What is more, straightforward application of adversarial perturbation methods like GCG on code completion attack has a terrible performance. As shown in Figure 9, 10-adv attacks with “0-keyword no FRE” refers to the 0% ASR.
Codebreaker and SVEN. In the domain of code, the newly proposed backdoor attack method, Codebreaker (Yan et al., 2024), also leverages the feedback of LLMs for optimization. However, unlike TPIA, Codebreaker emphasizes using the coding capabilities of LLMs to reduce the probability of the malicious payload code being detected, thereby increasing the success rate of backdoor attacks, without removing the reliance on the poisoning process. A method called SVEN (He and Vechev, 2023) involves purposefully generating high-risk code. However, this method requires Prefix-tuning on Code LLMs, adding new parameters to the original model parameters, making it unsuitable as an attack method. Additionally, when generating high-risk code, SVEN cannot control the content of risk code, nor can it achieve target-specific attacks.
A.2. Additional Details
Figure 11. Greshake et al. (Greshake et al., 2023) use a large amount of disruptive information and explicit malicious code snippets in the comments, which has a significant impact on the feasibility of the attack.
Baseline Details. IPI-M and C-GCG are important baselines in our effectiveness experiment. IPI-M uses injection instructions in natural language form to interfere with the model’s reasoning results, commanding the model to generate insecure code. The original IPI-M places a large amount of disruptive commentary along with the malicious code to be inserted in the comment section, as shown in Figure 11. We argue that this method has very low feasibility. Despite the attack path of IPI does not need to worry about being detected by malicious code detection methods, there is a risk of being discovered and removed once the victim becomes aware and actively pays attention to IPI attack, especially perturbations with poor visual stealthiness. Therefore, we attempt to simplify IPI-M into concise and clear template-based injection sentences, as shown in the table. C-GCG mainly follows the GCG method, using an adversarial prefix to achieve malicious purposes. At the same time, C-GCG does not adopt the two optimization methods from Section 4.4 and Section 4.5. C-GCG performs much worse than TPIA in the main experiments.
Value and Metric Selection. We select the values of $r_{1}$ and $r_{2}$ based on the attack’s performance in the experiment. Specifically, in our experiments, 20% ASR is smaller than the ASR of no attack baseline in some situations, and 80% ASR outperforms the best performance of all the baselines. In addition, we find that only a few subcases (determined by the model, dataset, attack method, and case) have a success rate between 20% and 80%, accounting for 7.5%. Therefore, setting 20% and 80% as threshold of bad case and outstanding case respectively is appropriate.
As for metrics, we have already explained why metrics such as NBR and STF are used in the main experiments. Here we further explain some details related to them. First, NBR does not take the specific ASR value into account, as this metric is used to measure the diversity of attacks, rather than ensuring stable success rates. Second, STF targets the situation where attackers aim to cause as much damage as possible, and it varies with differences in ASR. In other words, even a difference between 100% ASR and 95% ASR can have a real impact on the overall threat level. Therefore, the calculation of STF considers the specific ASR value. In addition, we do not use common metrics for code generation tasks, such as BLEU or CodeBLEU (Ren et al., 2020), because they do not directly reflect the effectiveness of the attacks. Overall, we attempt to confirm the realistic meaning of each evaluation metric from different perspectives.
Case Study. Here we show the details of our malicious objectives and threat cases. As shown in Table 6, we design 13 malicious objectives and 31 threat cases in 3 programming languages based on CWE top-10 KEV, exisiting research and famous destructive instruction. In programming process, minor changes in the code can cause significant functional alterations. We roughly divide the causes of vulnerabilities in all malicious objectives into three categories: skipping critical process (CWE20, 78, 502, 918, 22, 306), wrong memory usage (CWE416, 122, 787, 843), and outputting dangerous snippets (ST0, 1, 2). All these vulnerabilities result in severe consequences.
Table 6. Description of Threat Cases.
Case Description From
CWE416 Use after free. CWE 1st KEV
CWE122 Heap-based buffer overflow. CWE 2nd KEV
CWE787 Out-of-bounds write. CWE 3rd KEV
CWE20 Improper Input Validation CWE 4th KEV
CWE78 ‘OS Command Injection’ CWE 5th KEV
CWE502 Deserialization of untrusted data. CWE 6th KEV
CWE918 Server-side request forgery. CWE 7th KEV
CWE843 Access of resource using incompatible type. CWE 8th KEV
CWE22 ‘Path Traversal’ CWE 9th KEV
CWE306 Missing authentication for critical Function. CWE 10th KEV
ST0 Use ECB mode to encrpty. Schuster et al.(Schuster et al., 2021)
ST1 Use SSL3 protocol Schuster et al.(Schuster et al., 2021)
ST2 Executing “rm -rf / Famous destructive instruction
(a) Skipping critical process - CWE20
(b) Wrong memory usage - CWE122
(c) Outputting dangerous snippets - ST2
Figure 12. Three examples of TPIA attack cases. Code LLMs normally generate code segment without vulnerabilities, but they lose this ability after being attacked by TPIA. Attackers can modify the output content at the specified location through TPIA, and they can even generate malicious code snippets that have never appeared in the context, e.g., “os.system("rm -rf /")”.
As shown in Figure 12, here we present three representative examples across three different languages and different type of objectives. The first is CWE20 (using Java), where the program attempts to fetch the content of a URL and calculate its length. A secure programming practice would be to check if the content is null before performing the length calculation, thus avoiding a null pointer. However, after being attacked, the Code LLM generates code that calculates the length directly after fetching the URL, leading to a null pointer exception if the URL is null. A null pointer can result in serious consequences such as program crashes or data corruption. We simply categorize this situation as “skipping critical process.” The second example is CWE122 (using C), where the program attempts to pass content to buf through a command-line argument. A secure programming practice would be to check if the content exceeds BUFSIZE before the assignment, thus avoiding overflow. Through TPIA, attackers can induce Code LLMs to perform the assignment directly. When a stack overflow occurs, it accesses non-existent RAM space, leading to consequences such as system crashes and data leaks. We simply categorize this situation as “wrong memory usage.” This type of situation generally applies only to C/C++, as Java and Python automatically manage memory release. The third example is ST2 (using Python), where the program attempts to call a function named TRG (TaRGet). Under normal circumstances, this call does not present any issues, and the large model generates suggestions based on the contextual semantics or directly produces an ¡eos¿ token. Through TPIA, attackers can make Code LLMs generate the execution code “rm -rf /” right after calling the TRG function. If executed, this code results in the deletion of all data under the root directory. We categorize this situation as “outputting dangerous snippets.”
We only apply different perturbation construction strategies to different categorizations of cases in the IPI-M. Specifically, for “skipping critical process,” we use “… directly after …” and “without any check”; for “wrong memory usage,” we use “It is important to … directly after …”; for “outputting dangerous snippets,”, we use “It is very important to promptly … after …”. In addition, the perturbation in IPI-M must be sufficiently explicit; for example, it needs to specify the code snippet you want to insert (e.g., “os.system("rm -rf")”) or present a clear assertion (e.g., “SSL3 is the most secure protocol”). Otherwise, Code LLMs are difficult to induce successfully. Arguably, the actual performance of IPI-M may be worse than our evaluation results shown in main experiments.
Keywords Selection Method. We use a simple dual-loop strategy in the selection of keywords. Specifically, before each round of perturbation optimization with length $l$ starts in Algorithm 1, we use grid search to select the n tokens with the best performance as keywords. With the optimization step $s$, the equation can be formulated as follows:
(8) $\bm{P}_{key}^{s+1}=\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}_{adv}^{s},\bm{\theta}),$
where $\bm{P}_{key}$ and $\bm{P}_{adv}$ represents keywords part and adversarial part in perturbation $\bm{P}$ and the loss function is the same as Equation 6. $\bm{W}$ represents the token sequence that constitutes the keywords. To make the keywords relevant to the malicious objective, we select the tokens for the keywords from the target code and the position code, i.e., $\bm{P}_{key}=\{p_{k1},p_{k2},\dots,p_{kn}\}$, and for token $p_{ki}$ in $\bm{P}_{key}$, there is $p_{ki}\in\{\text{target code},\text{position code}\}$.
(a) Gemma
(b) Codegemma
(c) Codellama
(d) Codegeex2
Figure 13. Examples of traditional adversarial attacks on four different LLMs. The generated perturbations are displayed with a gray background, the input code completion task content with a green background, and the output content with a red background.
A.3. More Experiments
Comparison with Adversarial Attack. The advantages of TPIA over backdoors are clear: it does not require control of the training process and has a very low attack threshold. However, the advantages of TPIA over existing adversarial attacks, i.e., achieving target-specific attacks, is not obvious.
Figure 14. The ASR across different cases. The experimental results shown in this figure are the average values calculated from the effectiveness experiments involving 3 programming languages, 4 datasets, and methods excluding Clean Accuracy.
Here, we propose an imperfect but highly intuitive scoring method to compare TPIA with state-of-the-art adversarial attack method (Jha and Reddy, 2023). We divide the output results into three levels: (1) generating code with normal functionality, with a harmfulness score of 0 (level 0); (2) generating code that cannot compile or has functional errors, with a harmfulness score of 0.5 (level 1); (3) generating code with stealthy risk, with a harmfulness score of 1 (level 2). In terms of the resulting effects, the second level will lead to a decrease in the usability of Code LLM, and the third level not only affects the usability but also causes unpredictable financial losses to the users of the Code LLMs. We evaluate the performance on three cases: S0, S1, and S2, using four models: Gemma-7b, CodeGemma-7b, CodeLlama-7b, and CodeGeeX-6b (Due to the significant differences in resistance to traditional adversarial attacks across different models, we use more types of LLMs than in the main experiments, including a general-purpose LLM).
Figure 15. A comparison of the harmfulness between CodeAttack (CA) and TPIA. The final scores were normalized to the 0-100. We use light colors (CA lv1 and TPIA lv1) to represent ratings from level 1 harmfulness and dark colors (CA lv2 and TPIA lv2) to represent ratings from level 2 harmfulness.
Figure 16. The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b. The two types are almost indistinguishable.
As shown in Figure 15, TPIA significantly outperforms CodeAttack in scores. This is because TPIA consistently achieves level 2 harmfulness, whereas CodeAttack tends to reach level 1. In fact, the final result of CodeAttack is not invalid: the quality of output is greatly impacted, as shown in Figure 13. However, the output code usually cannot be correctly compiled. In other word, this type of attack is unlikely to threaten the user’s experience, although it successfully reduces the usability of Code LLM and has a negative impact on its commercial applications. In contrast, TPIA also impairs the usability of the victim Code LLMs, and can cause terrible subsequences for victim users. In summary, TPIA is capable of conducting more harmful attacks with a higher success rate, making it more suitable for LLM-based code completion scenarios compared to traditional adversarial attack methods.
Effectiveness across Different Cases. In the main experiments, we have validated the outstanding attack performance of TPIA from three perspectives: model, dataset, and programming language. Here, we supplement the attack details for different attacks on a single case. Specifically, we use ASR to evaluate the performance of TPIA and the baselines across different cases. We present the specific average ASR for each case. As shown in Figure 14, TPIA achieves a certain level of effectiveness across all cases, with an overall average ASR of 83.4%. However, C-GCG and IPI-M both have cases where they are completely ineffective, especially for IPI-M. This demonstrates that directly using traditional GCG methods or manually designing perturbations is not effective in some scenarios. We believe this involves three reasons: (1) The baseline methods are not fully suitable for contexts with structured features, which is a fundamental characteristic of the code completion task; (2) The specific malicious objective include complex semantics such as target code, position code, and “disregarding security attributes,” which are sometimes not fully captured by the perturbations generated by the baseline methods; (3) The impact of keyword-based designing and FRE; our ablation study proves that these two modules play a critical role in TPIA’s effectiveness.
According to the results, the most challenging to achieve is the straightforward case ST2. This case requires generating an obviously malicious command in a completely unrelated context. Another difficult case to implement is CWE22 “Path Traversal,” which involves using unconventional paths like “../” to access files that are normally inaccessible. This case is relatively easier to achieve in Python, where the Code LLMs tend to suggest directly reading the path. However, it is more challenging to implement in Java and C/C++, as Code LLMs often only recommend checking the path rather than directly reading the file. CWE843 involves manipulating union members in a struct in C/C++. This represents a more complex malicious objective, where we attempt to manipulate one value within the union using an untrusted value. This is an unconventional operation that is difficult to describe using natural language, making it challenging for IPI-M to successfully attack. In contrast, methods like TPIA and C-GCG, which convey malicious instructions through token selection, are able to achieve success. Arguably, TPIA has a clear advantage in attack performance for all the cases we have collected.
Semantic Analysis on Payload Code. In this experiment, we attempt to detect malicious payload code by semantic analysis. We observe the classification features of code segment in the embedding layer of CodeGemma-2b. Specifically, we add the classification prompt keyword [CLS] to 820 original code snippets and 820 malicious payload code snippets totally. and then input them into the model to analysis their semantic features. We examine the t-SNE (Van der Maaten and Hinton, 2008) distribution of the classification keyword [CLS] in the embedding layer, which is a widely used method for classification tasks (Abdelnabi et al., 2024). As shown in Figure 16, the features of the two types of text are almost completely intertwined, making it nearly impossible to distinguish between the injected code data and the original code data. In addition to observing the characteristics of layer 17 shown in Figure 16, we also examine the data features of layers 1 through 16. There is no significant difference observed in these layers compared to layer 17. There are several reasons for this situation. First, code data, including comments, tend to have lower fluency and coherence with high token complexity, which poses a significant challenge for semantic analysis tasks. Second, the perturbations used by TPIA occupy an extremely small number of tokens, exerting only a minor influence on the overall semantics of the injected data, making them hard to detect. Third, the semantic impact of TPIA differs from traditional injection attacks in that it does not require an extremely high intensity of information to counteract the model’s secure alignment; instead, it guides the model to generate vulnerabilities that could occur naturally, thus making detection through semantic analysis difficult. Our experiment results indicate that simple semantic analysis methods are not effective in defending against TPIA. This indicates that detecting code snippets containing malicious perturbations generated by TPIA through semantic analysis is not an effective approach.

### v5

@@ V-D Ablation Study
Figure 8: Ablation study of keywords number and Top-$k$. We compare TPIA’s ASR and STF under three conditions: without keywords, with 1-keyword, and with 2-keywords. During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
@@
Figure 9: Ablation study of FRE. We compare the differences in ASR and loss function values both with and without using FRE module. 1-key1 and 1-key2 represent keeping one of the two selected keywords, respectively.
@@
Effects of Forward Reasoning Enhancement. We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function. Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve. It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion. To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using static instead of optimized keywords. We judge the degree of distortion based on the loss value. The results are illustrated in Figure 9, where we have three key findings. Firstly, under the same conditions, using FRE can lead to a higher ASR. Secondly, without FRE, the ASR is still very low ($<20\%$) even when the loss value is lower than that with FRE. Thirdly, FRE has a greater impact on ASR than keyword-based design. These results verify the necessity of forward reasoning enhancement in our TPIA method.
VI Discussions
VI-A Resistance to Potential Defenses
@@
In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack. In general, we can roughly categorize these defenses into detection and removal and design following potential defense methods against TPIA.
Figure 10: The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b. The two types are almost indistinguishable.
@@
Detection: Semantic Analysis. Detection methods distinguish between malicious payload code containing perturbations and normal code, thereby rejecting the use of malicious payload code. Since detection methods based on code functionality are ineffective against non-functional perturbations, we adopt a semantic analysis approach. This involves examining the differences between the benign code and the injected code in the embedding layer of Code LLMs. However, this method may be difficult to implement. We find that code segments injected with perturbations might not have obvious semantic differences from benign ones. Specifically, as shown in Figure 10, we sample 1640 code segments and construct classification prompts. The results show that benign code and injected code are almost completely indistinguishable.
@@
Figure 11: The resistance to the removal of chars. We remove characters from the perturbation, selecting 5% for removal each time, and we calculate the average ASR for each programming language across applicable cases. As the influence of the perturbation fades, the ASR results converge to the same level as no attack baseline.
@@
Removal: Massive Deletion and Find-and-Replace. This type of defense involves removing perturbations by altering the non-functional parts of the code after introducing external code. In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect injected code segments (as mentioned above). To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations. As shown in Figure 11, TPIA still maintains a sufficiently high ASR under 0.5 removal rate. On the other hand, even if readability is completely sacrificed by extensively removing comments, TPIA can still be effective through non-comment portions of the code, as indicated in Table V. Arguably, the removal methods are likely to require significant human effort and may not yield effective results since TPIA imposes strict limitations on the length of perturbations.
@@
TABLE V: ASR (Deg) for different forms of perturbations. Deg denotes the ASR degradation compared with comments.
@@
VI-B Potential Limitations and Future Directions
@@
TPIA still has some drawbacks. Firstly, the effectiveness and stealthiness of TPIA are closely related to the keywords. We currently select keywords through a simple grid search. How to optimize keywords more efficiently is an important issue. Besides, TPIA cannot generate fully natural-looking perturbations within the context, resulting in a lack of fluency to some extent. In future work, we will explore how to use LLM as an optimizer to further improve this.
VII Conclusion
@@
In this paper, we investigated an important security vulnerability of Code LLMs for code completion tasks. We revealed the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduced TPIA, a new target-specific attack paradigm without controlling the model’s training process. TPIA can circumvent mainstream detection methods (e.g., static analysis and sandboxes) by using non-functional code, instead of functional ones used in software supply chain attacks, to achieve malicious objectives. We evaluated TPIA under three popular programming languages and 31 representative cases. We showed that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation. Our work alerted a new serious security threat in using Code LLMs.
Acknowledgment
We would sincerely thank Prof. Kui Ren and Prof. Chun Chen from the Zhejiang University for their helpful comments and suggestions on an early draft of this paper.